Assignment 2:

(a) Does either country have an absolute advantage in the production of wheat or beef? Explain.

(b) What is the opportunity cost of wheat in each country?

(c) What is the opportunity cost of beef in each country?

(d) Analyze comparative advantage and opportunities for trade between Home and Foreign.

                  Wheat        Beef

Home         600             200

Foreign      100             600
Answer #1

a.

Absolute advantage:

This is the most efficient way of production by which higher production could be achieved.

H has absolute advantage in wheat, since 600 is higher than 100.

F has absolute advantage in beef, since 600 is higher than 200.

b.

Opportunity cost:

This is the benefit foregone because of producing wheat but not beef.

Table

Country

Opportunity cost of wheat

Wheat

Beef

Opportunity cost

Home

600

200

200/600 = 0.33

Foreign

100

600

600/100 = 1

c.

Opportunity cost:

This is the benefit foregone because of producing beef but not wheat.

Table

Country

Opportunity cost of beef

Wheat

Beef

Opportunity cost

Home

600

200

600/200 = 3

Foreign

100

600

100/600 = 0.17

d.

Comparative advantage:

The country has lower opportunity cost should enjoy comparative advantage.

H has comparative advantage in wheat, because 0.33 is smaller than 1.

F has comparative advantage in beef, because 0.17 is smaller than 3.

Opportunities:

The country has comparative advantage should export the good to other country.

H should export wheat to F.

F should export beef to H.